Scenario III Another student, this time a biology undergraduate student (interested in biophysics), decides to travel to Bethesda MD (NIH) to study the effects of electromagnetic radiation on normal skin cells in white mice. Over the course of the summer, he irradiates a portion of the mouse skin using IR radiation (6000 nm), gamma 15 radiation (10 Hz), visible radiation (600 nm), UV radiation (2 X 10 Hz), and microwave radiation (10 cm) By using electron microscopy and other techniques, he studies the damage to the normal cells. Surprisingly, to him, his results were not quite what he expected. He then excises a portion of the mouse skin, dries the samples in an oven at 110°c, and again irradiates the skin samples. This time, his results are as he expected. Determine either the frequency (in Hz) or the wavelength (in nm) for each of the electromagnetic radiation samples the student used during the summer Determine the energy (in J) of one photon of each of the electromagnetic radiations the student used during the summer Arrange the electromagnetic radiation in order of decreasing energy In the normal skin cells, the student noted that there were two different types of damage, one caused by the higher energy photons and one caused by the lower type of damage the student saw ergy photons. Predict what One type of damage led to skin lesions (melanomas) being formed. Which types of photon would cause a melanoma? Explain why? The other type of damage is due to a completely different reason and has to do with whether the skin cells were hydrated or dehydrated. Which photon of radiation would most likely cause this type of damage and why?
Data Some Standard Enthalpies of Formation at 25 C. AMAHkcal Bond Energy/kJ mol Boad Compound wmol Compound mol mol mol Br-Br C-C C-C 193 343 615 812 24181-6779 AgCs) AgN,(s) Ag,08) AlO.ja -127 068-30 35H,Oig) +620 6148 3H,00 -285 -68 3 CC C-Ci C-F 741H,0 -1675.-400 40H,Sig) 0.00HgOa) -31.0 -187 7844 86 326 490 416 -20 63 4.93 Br-0 -90 83 - 21.70 0.C 290 CN C-O C-o C O(in CO C O C-C F-F H-Br H-CI H-H H-F H-I H-N 891 35 Bzig) +30 907 +7385) 00 +14.92 436 747-104.36 393.798-94.097 -601.7-143.77r -11.02 00 Cls), gaphite 00 0.00 g) +62 438 724 799 1046 244 158 366 432 436 CO damend+1895 +0453 kCIa) -17.88)Bria) 110 525 41 MgO(s) 393 509-9405NH) +226 7354 18NOg) +52 26 +12 49NO,lak 84 6820 23N,Os) +49 0311.72Ng) -635 09-151.77NaBrs -1205 92-288 39NaClis) -37.59p,(s) CH() COig) Coig) -74.81 46.11 CHg) 90 25 21 57 +7 93 +33 18 568 298 CH 9.16 2.19 391 -29.80 361 062 626 -98.24 -124 H-S N-N N-N N-O O-H O-o O O0 (in O 367 160 946 176 464 144 498 Caoa) Caco () Cuola) Fe,0,(a) Herig HCHg Hig) 411.153 -157 +34.11 +142 824 2-196 990-ig) -364 -296 83 - 70 93 -395.72 -94.56 -348 20 -83 22 470S0_ig) -92 307-22 06Ol) +26 48 From FTWall, Chemical Thermodmis, ed od. w. H. Froeman San Francion, 1974p.63 46 33 a 27 4 e au as ae Ce m e 4 se e N Te Rs Ta La ThDy La h-6.63X 10 J] c = 3X10' ms" Av-c v- 620X10"X ) 1 he E-IN iv-c- mv Tem) Temy-s- AH modat
Answer #1

1). Wavelength

IR 10^3-10^5 nm

Gamma 10^-3-10nm

Visible 400-700 nm

UV 10-300nm

Microwave 10^9nm

2). Energy of a photon is calculated as

E=hv=hc/lambda

V= frequency

IR

Wavelength= 6000nm=6000×10^9m

E= 6.63×10^-34 ×3×10^8/6000×10^9=3.3×10^-38J

Gamma

V= 10^15

E= 6.63×10^-34×10^15=6.63×10^-19J

Visible

Wavelength=600nm

E= 6.63×10^-34×3×10^8/600×10^9=3.3×10^-37J

UV

Frequency= 2×10^2Hz

E= 6.63×10^-34×2×10^7=13.26×10^-27J

Microwave

Wavelength=10cm

E= 6.63×10^-34×3×10^8/0.1=19.89×10^-25J

3). Decrease energy order

Gamma>UV>Visible>IR>Microwave

4). UV radiation have less energy and are less penetrating causing skin cancer and another damage is by microwaves.

5). UV radiation cause skin lesions.

6). Microwave radiation cause dehydration.